Hubble update: Not returning to normal today after all
I missed today's press briefing on Hubble, but there's enough detail in the report posted to the NASA website for a status update:
Hubble resumed the science timeline at Noon ET on Thursday, October 16. The first activities out of that on-board science timeline were the commanding of the science instruments from their safe to operate modes. This occurred nominally for Wide Field Planetary Camera 2 and the Near Infrared Camera and Multi Object Spectrometer. However, an anomaly occurred during the last steps of the commanding to the Advanced Camera for Surveys. At 1:40 pm, when the low voltage power supply to the ACS Solar Blind Channel was commanded on, software running in a microprocessor in ACS detected an incorrect voltage level in the Solar Blind Channel and suspended ACS. Then at 5:14 pm, the Hubble spacecraft computer sensed the loss of a "keep alive" signal from the NASA Standard Spacecraft Computer in the Science Instrument Command and Data Handling system and correctly responded by safing the NSSC-I and the science instruments. It is not yet known if these two events were related.
The investigation into both anomalies is underway. All data has been collected and is being analyzed. The science instruments will remain in safe mode until the NSSC-I issue is resolved. All other subsystems on the spacecraft are performing nominally and astrometry observations [with the Fine Guidance Sensors] continue.